DCNS Residential Computing Consultants

Maintained by Michael Barrow <mlbarrow@mit.edu>

Dormitory  RCC Contact Name  Email  Phone Room
620 Huntington Francis Hsu  fihsu  541-3212 123
Baker  Holly Grabowski  hollyg  5-7306 306
Baker  Iyad Obeid  iobeid  5-7296 546
Bexley  Barrett Comisky  elwood  5-9647 407A
Burton  Brandon Porter  bwporter  5-8383 Burton 432D
Burton  Catherine Preston  cassie  5-8307 Burton 332C
East Campus  Jack Fu  brainiac  5-6474 Bemis 414
East Campus  Hasan Sharif  has  5-6249 Walcott 204
East Campus  Srikar Srinath  srikar  5-6351 Walcott 306
MacGregor  Daniel Stevenson  danspot  5-9106 J116
MacGregor  Rupert Young  rcbear  5-9410 E321
McCormick  Nayana Ghantiwala  nayana  5-8681 721
McCormick  Aafia Siddiqui  aafia  5-8493 233
New House  Michael Ford  cthomas  5-7541 3-505
New House  Christiana Toutet  doidoi  5-7656 6-206
Next House  Eric Choi  echoi  5-8732 232
Next House  Susan Kim  hykim  5-8906 256
Random Hall  Jakub Roth  jroth  5-7348 411
Senior House  Mark Chandler  markch  5-6209 R204E

The following people are not RCCs:

IS DCNS  Jessie Stickgold-Sarah sorokin
IS DCNS  Michael Barrow  mlbarrow
IS DCNS  DCNS RCC archive  dcns-rcc-mtg@menelaus.mit.edu
IS DCNS  David Wong  wong

Current contacts as of March 19, 1994.

DCNS ILG Network Contacts

Maintained by Michael Barrow <mlbarrow@mit.edu>

ILG Name Address  -  City ILG Contact Name  Email  Phone
ADP  351 Mass Ave  - Cambridge Jason Stewart  keir  576-2792,3
AEP  155165 Bay State Rd. -  Boston Ron Schmelzer  deluded  859-8024
AP  479 Comm Ave  -  Boston Nicole Wainwright  naw  x5-1510
ATO  405 Mem Dr  - Cambridge Mike Schlosser  schloss  252-6520
BTP  119120 Bay State Rd -  Boston Eric Mudama  edmudama  536-4269
CP  32 Hereford St  -  Boston Han Shu  shu  353-1795
DKE  403 Mem Dr  - Cambridge Andy Kass  delphi  494-8250
DP/No. 6 428 Mem Dr  - Cambridge Pekka Kauranen  pekka  577-0085
DTD  416 Beacon St  -  Boston Mark Hartge  hardguy  236-0987
DU  526 Beacon St  -  Boston Jon Litt  littlitt  536-3931
ET  259 St. Paul St  - Brookline Phillip Tiongson  sdallas  734-9211
Fenway  34 The Fenway  -  Boston Tom Yu  tlyu  437-1043
KS  407 Mem Dr  - Cambridge Owen Johnson  myriad  494-0330
LCA  99 Bay State Rd  -  Boston Brian Coan  bscoan  236-1159
ND  460 Beacon St  -  Boston Travis Ackins  travis  437-7300
PBE  400 Mem Dr  - Cambridge Hector Ayala  ayala  x3-2843
PDT  97 Bay State Rd  -  Boston Jon Michelson  jonm  247-8691
PGD/Fiji 28 The Fenway  -  Boston Robert D’Onofrio  radonofr  262-3529
PKS  530 Beacon St  -  Boston Matt Domsch  luigi  536-3683
PKT  229 Comm Ave  -  Boston Adam Taylor  ataylor  437-7795
PSK  487485 Comm Ave  -  Boston Jeff VanDyke  jvandyke  267-2199
PLP  450 Beacon St  -  Boston Alexander Rosen  arose  267-5270
pika  69 Chestnut St  - Cambridge Craig Henderson  krandor  492-6983
SAE  484480 Beacon St  -  Boston Jeremy Pitcock  pit  266-3575
SC  532 Beacon St  -  Boston Walter Koch  waltvk  262-3192
SPE  515 Beacon St  -  Boston Sean Hogarty  hogarty  536-1300
SH  111 Bay State Rd  -  Boston Soloman Douglas  srcd  247-0506
TEP  253 Comm Ave  -  Boston Bill Helfinstine  bhelf  262-5090
TC  528 Beacon St  -  Boston Darius Miller  dmiller  267-1801x132
TDC  372 Mem Dr  - Cambridge Marc Bejarano  beej  494-8300
TXI  6466 Bay State Rd  -  Boston Mike Evans  mike e  266-2827
WILG  355 Mass Ave  - Cambridge Hillary Hudis  hudidis  x3-6799
ZBT  58 Manchester Rd  - Brookline Daniel Bromberg  ddaniel  731-2685
ZP  233 Mass Ave  - Cambridge Bill Euerle  wjeuerle  661-4111

The following people are not ILG Network Contacts:

IS DCNS E40-331  Michael Barrow  mlbarrow  x3-7664
IS DCNS E40-342g  Mark Curby  mlc  x3-7725
CP  Corporation Advisor  tholtey@world.std.com
CP  Backup  hogjordn
CP  Backup  viny
Fenway  Advisor  Jesse Stickgold-Sarah sorokin
SH  Advisor  Dave Mycue  wolvie@otw.tch.harvard.edu
ZBT  Advisor  Jin Choi  jsc
WILG  Corporation Advisor  carney@buenga.bu.edu
WILG  Backup  whimsy

Current contacts as of March 19, 1994.